OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. -18/35 / 22/25

OpenStudy (anonymous):

See: \[\huge \color{green}{\frac{\frac{a}{b}}{\frac{c}{d}} = \frac{a \times d}{ b \times c}}\] Try to use this concept..

OpenStudy (anonymous):

-18/35 / 22/25.......a x b/b x c = 450-770

OpenStudy (anonymous):

I meant 450/770

OpenStudy (anonymous):

How?? can you show me properly?/ Where is negative sign gone??

OpenStudy (anonymous):

Can you solve for: \(\large \frac{-450}{770}\) by taking the factors common and then cancelling them.. Simply reduced it to simplest fraction..

OpenStudy (anonymous):

I have no idea what I did, ok I will try...Thanks!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (usukidoll):